OpenStudy (anonymous):

John and Martha each had a steak dinner last night at Fred's diner. The steaks were $12.95 apiece, and they each had a glass of wine at $6 apiece. The bill included an additional 8% tax. How much was the total of John and Martha's bill?

OpenStudy (carson889):

2 * $12.95 + 2 * 6 = 25.90 + 12 = 37.90 before additional tax. 37.90 multiplied by 8% is: 37.90 * 0.08 = 3.03. This the total bill was $40.93. We could also just multiplied the subtotal ($37.90) by 1.08 to get the final total, instead of finding the tax value and then adding it to the subtotal.